OpenStudy (anonymous):

(Averages: Means, Medians, and Modes) 1. What is the average (arithmetic mean) of 8, 7, 7, 5, 3, 2, and 2? A. 34/7 B. 45/6 C. 46/7 D. 5 E. 64/5

OpenStudy (mertsj):

Add them together and divide by 7

OpenStudy (anonymous):

I understand that the answer is 4 6/7 but how can I break it down to understand how I got the answer. I know that I have to sum all of the values and divide them by the number of values in order to get the answer but I'm not understanding how the 6/7 came into place?

OpenStudy (mertsj):

\[\frac{8+7+7+5+3+2+2}{7}=\frac{34}{7}=4 \frac{6}{7}\]

OpenStudy (anonymous):

I did add them and got 34 divided by 7=4

OpenStudy (mertsj):

34 divided by 7 is NOT 4

OpenStudy (mertsj):

So mark choice A and go on to the next problem.

OpenStudy (anonymous):

Ok, well the correct answer is C. which is 4 6/7

OpenStudy (mertsj):

They are the same thing.

OpenStudy (anonymous):

How did you get the 6?

OpenStudy (anonymous):

Oh ok. Thanks.

OpenStudy (mertsj):

How are we supposed to know that C is 4 6/7 and not 46/7 as it says?

OpenStudy (mertsj):

Does that mean that A is 3 4/7 or is it 34/7?

OpenStudy (mertsj):

When you post stuff it is very important that it be posted accurately.

OpenStudy (anonymous):

I have no idea? LOL! Just trying to figure out how to get the correct answer that's all. Thanks!

OpenStudy (anonymous):

It was posted the way I saw it.

OpenStudy (mertsj):

yw